# Recent quantum runtime (dis)advantages

### v2

@@ Abstract
We (re)evaluate recent claims of quantum advantage in annealing- and gate-based algorithms, testing whether reported speedups survive rigorous end-to-end runtime definitions and comparison against strong classical baselines. Conventional analyses often omit substantial overhead (readout, transpilation, thermalization, etc.) yielding biased assessments. While excluding seemingly not important parts of the simulation may seem reasonable, on most current quantum hardware a clean separation between “pure compute” and “overhead” cannot be experimentally justified. This may distort “supremacy” results. In contrast, for most classical hardware total time $\approx$ compute ++ a weakly varying constant leading to robust claims. We scrutinize two important milestones: (1) quantum annealing for approximate QUBO [PRL 134, 160601 (2025)], which uses a sensible time-to-$\varepsilon$ metric but proxies runtime by the annealing time (non-measurable); (2) a restricted Simon’s problem [PRX 15, 021082 (2025)], whose advantageous scaling in oracle calls is undisputed; yet, as we demonstrate, estimated runtime of the quantum experiment is $\sim 100\times$ slower than a tuned classical baseline. Finally, we show that recently claimed “runtime advantage” of the BF-DCQO hybrid algorithm (arXiv:2505.08663) does not withstand rigorous benchmarking. Therefore, we conclude that runtime-based supremacy remains elusive on NISQ hardware, and credible claims require a careful time accounting with a proper reference selections, and an adequate metric.
@@ I Introduction
The main motivation for the development of quantum computers, noted in pioneering works on the subject [1, 2], is the possibility of efficiently solving problems currently intractable for classical computation. Initially a theoretical field, quantum computing has seen rapid technological progress, and major roadmaps predict error-corrected machines by around $2030$ [3, 4, 5]. Annealing technology is also advancing [6]. Yet, demonstrating quantum advantage experimentally for both these paradigms remains a key challenge driving research.
@@
Quantum advantage can be defined in various ways, such as in computation or metrology [7]. One approach relies on complexity-theoretic arguments proving better scaling than classical algorithms, as in Simon’s algorithm [8], which shows exponential query separation, or Shor’s algorithm [9], based on the presumed hardness of factoring [10]. While generic speedups for NP-hard problems are unlikely, restricted instances may allow improvements [11], motivating studies of heuristic quantum algorithms [12]. These may provide complexity statements for specific cases or empirical advantages, given appropriate hardware. Other notions, such as energetic advantage, have also been discussed [13].
@@
As quantum devices improve, attention turns to experimental verification [14]. Early claims based on Random Circuit Sampling [15] were challenged by classical simulation [16, 17], though recent results seem intractable [18]. Further claims include favorable time-to-solution scaling for QUBO optimization on D-Wave annealers [19], query complexity advantage in restricted Simon’s problem [20], and runtime benefits for a hybrid classical-quantum algorithm [21]. These claims are significant as they reflect the growing technological maturity of quantum computing. However, a key question remains: do these reported advantages translate into actual runtime improvements — a critical factor for practical adoption? In this work, we reassess various strategies employed in these experiments and show that the answer is negative: no runtime advantage is observed when runtime is properly measured, and an appropriate classical baseline is selected.
In particular, in Sec. II, we examine the findings of Ref. [19], which reported that the runtime of a quantum annealing-based approximate QUBO scaled more favorably than the best classical reference algorithm (therein PT-ICM [22]). We discuss challenges in defining total runtime for such experiments and show that, under a robust definition, no advantage remains. Next, we examine the results of Ref. [21], where the digitized counter-diabatic quantum optimization algorithm, a gate-based approach, was reported to achieve a runtime advantage over classical Simulated Annealing (SA) [23] and CPLEX [24]. We identify several issues with the definitions of classical and quantum runtime used therein, which ultimately eliminates this advantage altogether.
@@
In Sec. III, we address the reported query complexity advantage for a restricted Simon’s problem [20] on noisy gate-based IBM quantum computers. We investigate whether fewer oracle calls lead to shorter runtimes and find that, although the classical algorithm indeed exhibits exponentially worse scaling in oracle calls, in the regime considered in [20] its runtime is actually two orders of magnitude shorter than the quantum algorithm.
@@
Finally, in Sec. IV, we discuss the importance of selecting an appropriate classical reference algorithm and, focusing on runtime as the key metric, outline essential criteria for this choice. We revisit the reference selection made in [21] and argue that a proper choice already eliminates any quantum-classical runtime advantage. This scenario closely parallels a recent discussion in [25], where it was shown that an appropriately chosen classical baseline can demonstrate robust scaling and effectively remove claims of quantum supremacy in discrete approximate optimization (cf. Ref. [19]) under operationally meaningful conditions. For completeness, we also verified the performance of a D-Wave quantum annealer on the problems studied in [21], and found that adiabatic quantum computing offers no runtime advantage either. On the contrary, these instances appear to be particularly challenging for the D-Wave machine, as they require a complex embedding that ultimately degrades solution quality.
@@ II Quantum advantage in algorithm runtime
where $J_{ij}$ are coupling strengths between spins, and $h_{i}$ are local magnetic fields [27].
@@
where $t_{f}$ is the time taken to generate the solution and $p_{E\leq E_{0}+\varepsilon|E_{0}|}$ is the probability of finding a solution with energy $E$, within $\varepsilon$ optimality gap of the true ground state energy $E_{0}$ in cases when the solution is known, or a reference energy otherwise (e.g. the best energy found by a reference solver). In most cases the probability distribution $p_{E\leq E_{0}+\varepsilon|E_{0}|}$ remains unknown, and it needs to be replaced by an estimator, whose computation involves firstly a non-trivial optimization of the solver settings, and subsequently a number of independent runs of the solver with optimal parameters, from which the estimate is established. The employed function of success probability estimates the number of times an algorithm needs to be run to find a solution within the desired optimality.
Investigation of $\mathrm{TT}\varepsilon$ for specific classes of optimization problems, or its scaling with the problem size serves as a standard comparison tool between different heuristic algorithms, including the quantum ones [19]. However, it is clear from the above definition that a proper definition of the runtime $t_{f}$ is absolutely crucial for fair and unbiased comparison.
In most cases defining and measuring runtime for classical algorithms poses no problem. The set of tools for straightforward runtime measurements is also available [29, 30, 31]. On the other hand, there are two fundamental challenges concerning quantum runtime. The first one concerns the proper definition of quantum runtime, as one must determine which of various quantum computation stages should be included in the total runtime. Secondly, one needs to establish a robust method of measuring times of those stages. In the subsections these issues would be addressed separately for quantum annealing and gate based devices.
II.1 Runtime for annealing quantum devices
Solution of an optimization problem on a quantum annealer consists of the following stages. Firstly, an optimization problem must be loaded into the annealing device. This involves two steps: embedding, which maps a given instance of the Ising model onto the device topology, and programming, which sets the physical properties of the device to realize the embedded problem. Subsequently, the quantum annealing protocol takes place. Here it is important to note that the duration of the annealing is an input parameter to the protocol, and, at least using the standard cloud access, it is not possible to measure it independently [32]. Due to the probabilistic nature of the annealing protocol, multiple runs are required to increase the probability of obtaining a high-quality solution. These annealing runs are executed sequentially, and are interrupted by the two processes. The first is readout, which allows to obtain information about the final state of each protocol, whereas the second is thermalization, which restores the device to its initial state. Both processes have their intrinsic runtimes.
It seems reasonable, therefore, that for quantum annealers, the total runtime should be defined as the sum of all annealing runs, including readout, thermalization, and programming times. This is not a common practice; for instance, in [19], runtime was defined (and not measured) solely as the annealing time, excluding other contributions. Our aim is now to examine how this decision influenced the reported scaling of $\mathrm{TT}\varepsilon$, and how the scaling changes when the additional stages of the annealing protocol are taken into account. To this end we repeated the experiments using the same instances as in Refs. [19, 33]. When $t_{f}$ is set to the annealing time per sample, we reproduce the results of [19], confirming that the methodology was implemented correctly. We then investigated how the $\mathrm{TT}\varepsilon$ scaling changes when $t_{f}$ is taken as the total QPU time, which includes programming, annealing, readout, and thermalization times, as reported by the cloud interface [34]. To further validate these results, we measured the usage time of the cloud interface itself, which also accounts for access overheads.
In both cases, the results indicate that the runtime is almost constant. The analysis of the scaling exponent support this findings as the uncertainties of the fits are too large to conclude that the exponents are non-zero. Analysis of the different runtime components shows that the dominant contribution to the total runtime comes from the readout, which is on the order of $200\mu s$ per annealing run. This implies that measuring the final quantum state of the D-Wave quantum annealer is up to two orders of magnitude longer than executing the quantum evolution required to prepare this state, as the annealing time in this experiment varies from $0.5\mu s$ to $27\mu s$. This discrepancy, and the difficulties it poses for investigating $\mathrm{TT}\varepsilon$ scaling, motivated the exclusion of readout time in [19]. However, this argument is not entirely convincing. Measurements are an inescapable part of quantum computing, whether analogue or digital. Information encoded in a quantum state must be read from a quantum device, as only the measurement results provide the solution to a computational problem.
@@ II.1 Runtime for annealing quantum devices
For reference we compared scaling of the quantum annealing protocol to a classical algorithm, Simulated Bifurcation Machine (SBM [35, 36, 25]), cf. App. A1. Since it is a classical algorithm, the runtime can be measured directly and split into components, in particular the pure computation time, in our case performed on GPU – $t_{f}^{\rm GPU}$, and the overhead time $t_{f}^{\rm overhead}$, which is dominated by the data transfer between CPU and GPU and hyperparameter tuning as exaplained in Appendix A1. We denote the total time from the end user perspective as $t_{f}^{\rm tot}=t_{f}^{\rm GPU}+t_{f}^{\rm overhead}$. In Fig. II.1 we show that the impact of overhead on the scaling of $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ is certainly non-negligible, albeit much less pronounced than in the quantum case. Our experiments show that runtime $\approx$ “compute” ++ “a weakly varying constant” does not hold for quantum annealers. In contrast, this approximation remains valid for classical hardware such as GPUs, which explains why scaling behavior is preserved across different methodologies for measuring time.
FIG. 1. : Time-to-epsilon $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ scaling with the size $N$ of the Sidon-28 instances, for values of $\varepsilon=0.75\%$ in panel (a), and $\varepsilon=1.25\%$ in panel (b). Results for QAC (blue) solver are reproduced from Ref. [19], courtesy of the authors. Remaining QAC data concerns experiments performed by us on the instances form Ref. [19]. With the same definition of runtime the results (orange) are consistent with  [19], what confirms correctness of methodology implementation. In addition $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ defined using the complete QPU access time $t_{f}^{{\rm QPU\,access}}$ as reported by D-Wave’s cloud interface (green), as well as runtime measured with cloud access $t_{f}^{{\rm runtime}}$ (red) are presented. Solid lines are power-law fits $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}\propto N^{\alpha}$, with the corresponding exponents $\alpha$ shown on the plot. The instance size range spanned by the lines denote which data points were used for the respective fits. Both plots clearly demonstrate that with the proper runtime definition $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ is constant for considered problem sizes. For comparison data for Simulated Bifurcation Machine (SBM) solver[36, 35, 37, 25] with $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ computed using 1 GPU and total runtime $t_{f}^{\rm tot}$ (cyan), as well as pure GPU runtime $t_{f}^{\rm GPU}$ (magenta) are presented. In this case the scaling of $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ is more robust with respect to different runtime definitions, which is generally the case for classical algorithms. The bottom panel (c) shows detailed data of the fitting exponent and its uncertainty for values of $\varepsilon\in\{0.5,0.75,1.00,1.10,1.25,1.5\}\%$
To properly address this issue, all stages of computation would need to be included. On such time scales, however, the runtime becomes nearly constant for the problem sizes considered. Conclusive scaling results for current annealing devices would therefore require experiments with annealing times exceeding $200\mu\text{s}$ and with substantially more than $5000$ qubits.
@@ II.2 Runtime for digital quantum devices
For digital quantum computers, one can distinguish computation stages analogous to those discussed for quantum annealers. The first stage is preprocessing, which involves translating a quantum circuit into the topology and native gate set of a chosen device, a process known as transpilation. The transpiled circuit is then executed multiple times (shots), with consecutive runs separated by readout and thermalization. A proper definition of runtime for digital quantum computers should therefore account for all these processes. This perspective was also advocated in [38], which describes a procedure for measuring their durations on IBM devices.
@@
With this definition of quantum runtime, we revisit the results of Ref. [21]. Their digitized counter-diabatic quantum optimization algorithm was reported to exhibit a runtime advantage over Simulated Annealing (SA [23]) and IBM’s proprietary CPLEX solver [24] for a class of Higher-Order Unconstrained Binary Optimization (HUBO) problems. These are probles whose cost function includes multi-variable terms of order $N=3$, thereby generalizing the Ising model of Eq. (1),
[…]
The choice of these two particular classical references is discussed in Sec. IV. Here, we focus solely on the employed definition of runtime, as well as its measurement. The algorithm considered in [21] is a hybrid quantum-classical algorithm consisting of two runs of SA and a single run of the quantum routine. This methods begins with a single run of SA providing an initial guess for quantum routine. The state produced by the quantum routine is then fed into a second run of SA, which aims to mitigate possible readout errors.
The total runtime thus comprises both classical and quantum components. The classical runtime was estimated based on the time required to perform a single pass of SA. Therefore, the reported times are not the result of a physical measurement, but rather estimates derived from the number of passes performed. Curiously, the authors did not account for overheads associated with SA, which, if included, would affect their conclusions, as the overheads provided by the authors are on the order of $1.6\mathrm{s}$. We emphasize that a robust protocol for measuring classical computational time would allow for avoiding such ambiguities. The quantum runtime was estimated under the assumption that the quantum devices used in the experiments can execute $10^{4}$ circuit repetitions per second. This estimate, however, excludes overheads such as programming and transpilation time, as discussed above. For such experiments a better methodology, specifically designed for IBM quantum devices, could be applied, cf. Ref. [38].
To conclude, there are a number of severe problems with the definition and measurement of runtime for both classical and quantum algorithms, which constitute the digitized counter-diabatic quantum optimization algorithm. Due to these concerns, and the lack of an appropriate metric, we conclude that at the present stage the runtime advantage claimed in Ref. [21] is unfounded. Furthermore, even if such a non-standard approach to measuring runtime were to be justified, the claimed “supremacy” would not withstand comparison against a strong classical reference. We explicitly address this point in Sec. IV, where we highlight further methodological problems in Ref. [21]. For instance, applying their methodology to the same class of problem instances shows that other classical methods (and even optimized SA) can achieve comparable or superior results, thereby undermining the BF-DCQO supremacy claim. A detailed analysis is provided in App. A3.

### v3

@@ Abstract
A robust definition of quantum runtime is essential for assessing the performance of quantum algorithms and claims of quantum advantage. While for most classical hardware the total runtime is well approximated by computation plus a weakly varying constant, on current quantum hardware a clean experimental separation between “pure computation” and “overhead” is often not justified. Consequently, conventional quantum runtime analyses that exclude substantial system-level overheads (e.g., readout, transpilation, thermalization) can lead to biased performance assessments. In this work we introduce experimentally grounded, end-to-end definitions of quantum runtime for both digital and analogue quantum computers, together with a methodology for selecting strong classical baselines for quantum–classical runtime comparisons. Within this framework, we evaluate recent claims of quantum advantage in annealing- and gate-based algorithms. We examine three representative case studies. First, we revisit quantum annealing for approximate QUBO problems [Phys. Rev. Lett. 134, 160601 (2025)], which employs a well-motivated time-to-$\varepsilon$ metric but effectively uses annealing time as a proxy for runtime. Second, we analyze a restricted implementation of Simon’s problem [Phys. Rev. X 15, 021082 (2025)], where the favorable scaling in oracle calls is undisputed; however, we show that the estimated wall-clock runtime of the quantum experiment is approximately two orders of magnitude slower than a tuned classical baseline at the tested sizes. Finally, we find that the recently reported runtime advantage of the BF-DCQO hybrid algorithm (arXiv:2505.08663) is not observed under more comprehensive benchmarking. Therefore, on current NISQ hardware, runtime-based quantum advantage has not yet been demonstrated under experimentally grounded performance metrics, and credible claims require careful time accounting, appropriate performance measures, and properly chosen classical reference implementations, as discussed in this work.
@@ I Introduction
The main motivation for the development of quantum computers, as noted in pioneering works on the subject [1, 2], is the possibility of efficiently solving problems that are currently intractable for classical computation. Initially a theoretical field, quantum computing has seen rapid technological progress, and major roadmaps now predict error-corrected machines by around $2030$ [3, 4, 5]. Quantum annealing technology is also advancing [6]. Despite this progress, experimentally demonstrating quantum advantage for both paradigms remains a central challenge, driving ongoing research [7].
@@
Quantum advantage can be defined in various ways, including advantages in computation or metrology [8]. One prominent approach relies on complexity-theoretic arguments that establish more favorable scaling than classical algorithms, as exemplified by Simon’s algorithm [9], which exhibits an exponential separation in query complexity, or Shor’s algorithm [10], based on the presumed hardness of factoring [11]. While generic speedups for NP-hard problems are widely believed to be unlikely, restricted instances may admit improvements [12], motivating studies of heuristic quantum algorithms [13]. Such approaches may provide either complexity-theoretic guarantees for specific problem classes or empirical performance improvements on suitable hardware. Other notions of advantage, such as energy efficiency, have also been discussed [14].
@@
As quantum devices continue to improve, increasing attention has turned toward experimental verification of quantum advantage claims [15]. Early demonstrations based on random circuit sampling [16] were subsequently challenged by classical simulation techniques [17, 18], although more recent results appear to lie beyond the reach of known classical methods [19]. Additional experimental claims include favorable time-to-solution scaling for QUBO optimization on D-Wave annealers [20], query-complexity advantages in restricted implementations of Simon’s problem [21], and reported runtime benefits for hybrid classical–quantum algorithms [22]. These studies reflect the growing technological maturity of quantum hardware and are therefore of considerable interest. At the same time, they raise an important question: Do quantum performance assessment methodologies currently in use – particularly those based on scaling arguments or proxy metrics – result in robust quantum advantage with actual end-to-end runtime improvements critical for practical adoption?
Specifically, in Sec. II, we discuss challenges associated with defining total runtime in annealing- and gate–based experiments, and propose consistent and operationally meaningful quantum runtime definitions for both quantum computing paradigms. Using those definitions we revisit two recent results regarding quantum advantage for analogue and digital quantum algorithms.
Firstly we address Ref. [20], which reported favorable scaling of a quantum annealing–based approximate QUBO solver relative to a classical reference algorithm (PT-ICM [23]), and show that, when a rigorous end-to-end runtime definition is employed, the reported advantage is no longer observed.
Next, we analyze the results of Ref. [22], where a digitized counter-diabatic quantum optimization algorithm – a gate-based approach – was reported to achieve a runtime advantage over classical simulated annealing (SA) [24] and CPLEX [25]. We find that also in this case the use of the comprehensive runtime definition does not allow concluding that the considered algorithm provides a runtime advantage.
@@
In Sec. III, we address the reported query-complexity advantage for a restricted implementation of Simon’s problem [21] on noisy gate-based IBM quantum processors. We examine whether reduced oracle-call complexity leads to shorter wall-clock runtimes and find that, while the classical algorithm exhibits exponentially worse scaling in oracle calls, its actual runtime in the regime explored in Ref. [21] remains approximately two orders of magnitude shorter than that of the quantum implementation.
@@
Finally, in Sec. IV, we discuss the importance of selecting appropriate classical reference algorithms when evaluating runtime performance. Focusing on runtime as the primary metric, we outline key criteria for classical baseline selection and revisit the choices made in Ref. [22]. We show that adopting an alternative, well-motivated classical reference is sufficient to remove the reported runtime advantage. This situation parallels recent discussions in Ref. [26], where the use of stronger classical baselines was shown to alter conclusions regarding quantum advantage in discrete approximate optimization problems (cf. Ref. [20]) under operationally meaningful conditions. For completeness, we also evaluate the performance of a D-Wave quantum annealer on the instances studied in Ref. [22] and find that, in this setting, adiabatic quantum computing does not yield a runtime advantage. Instead, these instances prove particularly challenging for the annealer, as the required embedding significantly degrades solution quality.
@@ II Quantum advantage in algorithm runtime
where $J_{ij}$ are coupling strengths between spins and $h_{i}$ are local magnetic fields [28].
In the case of approximate optimization, a standard figure of merit is the time-to-epsilon, $\mathrm{TT}\varepsilon$ (sometimes used interchangeably with the time-to-approximation-ratio, $\rm TT$$_{\mathcal{R}}$ [29]), which quantifies the time required to obtain a solution whose energy lies within an $\varepsilon$ fraction of the ground-state energy. This quantity is defined as
where $t_{f}$ denotes the time taken to generate a single solution and $p_{E\leq E_{0}+\varepsilon|E_{0}|}$ is the probability of obtaining a solution with energy $E$ within an $\varepsilon$ optimality gap of the true ground-state energy $E_{0}$, when this value is known, or relative to a reference energy otherwise (e.g., the best energy found by a reference solver). In most practical settings, the probability distribution $p_{E\leq E_{0}+\varepsilon|E_{0}|}$ is unknown and must be estimated. This typically involves non-trivial optimization of solver parameters, followed by multiple independent runs using the optimized settings, from which the success probability is inferred. Its functional dependence then determines the expected number of repetitions required to obtain a solution within the desired optimality threshold.
Investigations of $\mathrm{TT}\varepsilon$ for specific classes of optimization problems, as well as studies of its scaling with problem size, are commonly used as comparison tools between different heuristic algorithms, including quantum approaches [20]. However, it follows directly from the above definition that a careful and consistent definition of the runtime $t_{f}$ is essential for fair and unbiased comparisons.
In most cases, defining and measuring runtime for classical algorithms is relatively straightforward, and a variety of mature profiling tools are available for this purpose [30, 31, 32]. In contrast, defining and measuring runtime for quantum algorithms presents two key challenges. The first concerns the operational definition of quantum runtime, namely which stages of the quantum computation pipeline should be included in the total runtime. The second concerns the practical measurement of the durations of these stages in an experimental setting. In the following subsections, we address these issues separately for quantum annealing and gate-based quantum devices.
II.1 Runtime for annealing quantum devices
Solution of an optimization problem on a quantum annealer consists of several stages. First, the optimization problem must be loaded onto the annealing device. This involves two steps: embedding, which maps a given instance of the Ising model onto the device topology, and programming, which configures the physical parameters of the device to implement the embedded problem. Subsequently, the quantum annealing protocol is executed. It is important to note that the annealing duration is an input parameter to the protocol and, under standard cloud access, cannot be independently measured by the user [33]. Due to the probabilistic nature of the annealing process, multiple runs are typically required to increase the probability of obtaining a high-quality solution. These runs are executed sequentially and are separated by two additional processes: readout, which retrieves information about the final state of each run, and thermalization, which restores the device to its initial state. Both processes contribute non-negligible time overheads.
The comprehensive total runtime definition of a quantum annealer consists of the sum of all annealing runs together with the associated readout, thermalization, and programming times. This definition is not universally adopted; for example, in Ref. [20], runtime was defined solely in terms of the annealing time, with other contributions excluded. Our goal here is to examine how robust is this choice of runtime definition in the context of $\mathrm{TT}\varepsilon$, and to what extent the scaling is affected when additional stages of the annealing protocol are incorporated.
To this end, we repeated the experiments using the same problem instances as in Refs. [20, 34]. When $t_{f}$ is identified with the annealing time per sample, we reproduce the results reported in Ref. [20], confirming that the methodology was implemented consistently. We then examine how the scaling of $\mathrm{TT}\varepsilon$ changes when $t_{f}$ is defined as the total QPU access time, including programming, annealing, readout, and thermalization, as reported by the cloud interface [35]. To further validate these results, we also measure the total cloud-access time, which includes additional access-related overheads.
@@ II.1 Runtime for annealing quantum devices
For comparison, we evaluate the scaling behavior of the quantum annealing protocol against a classical Simulated Bifurcation Machine (SBM [36, 37, 26]); see App. A1. As a classical algorithm, SBM allows direct measurement and decomposition of runtime into components, including the pure computation time, performed on a GPU ($t_{f}^{\rm GPU}$), and overhead time ($t_{f}^{\rm overhead}$), dominated by data transfer between CPU and GPU and hyperparameter tuning, as described in App. A1. We denote the total runtime from an end-user perspective as $t_{f}^{\rm tot}=t_{f}^{\rm GPU}+t_{f}^{\rm overhead}$. As shown in Fig. II.2, the impact of overhead on the scaling of $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ is non-negligible, though substantially less pronounced than in the quantum annealing case.
These results indicate that the approximation runtime $\approx$ “compute” ++ “weakly varying constant” does not hold for current quantum annealers, whereas it remains a reasonable approximation for classical hardware such as GPUs. This difference helps explain why scaling behavior for classical algorithms is typically more robust with respect to the choice of runtime definition.
[…]
Our results also allow us to identify another possible route toward a scaling quantum annealing advantage. This concerns a hypothetical scenario in which annealing technology progresses such that the number of qubits increases, while other relevant timescales – in particular the annealing and readout times – remain constant or grow only slowly. Under this assumption, and assuming no comparable improvements in the classical reference technology, a rough extrapolation of the data presented in Fig. II.2 suggests that a crossover between the SBM and annealing performance could occur for problem sizes on the order of $10^{5}$ variables.
@@ II.2 Runtime for digital quantum devices
For digital quantum computers, one can identify computation stages analogous to those discussed for quantum annealers. The initial stage involves preprocessing, in which a quantum circuit is translated into the topology and native gate set of a given device through transpilation. The transpiled circuit is then executed multiple times (shots), with consecutive runs separated by readout and thermalization steps. A comprehensive definition of runtime for digital quantum devices should therefore include all of these processes. This perspective is also adopted in Ref. [38], which outlines procedures for measuring these contributions on IBM quantum devices.
@@
Using this definition of runtime, we revisit the results of Ref. [22]. In that work, a digitized counter-diabatic quantum optimization algorithm was reported to achieve a runtime advantage over Simulated Annealing (SA [24]) and IBM’s proprietary CPLEX solver [25] for a class of Higher-Order Unconstrained Binary Optimization (HUBO) problems. These problems involve cost functions with multi-variable terms of order $N=3$, generalizing the Ising model in Eq. (1),
$P(s)=\sum_{\begin{subarray}{c}i_{1},\dots,i_{N}\\ i_{1}+\cdots+i_{N}\leq 3\end{subarray}}J_{i_{1}\dots i_{N}}s_{1}^{i_{1}}\dots s_{N}^{i_{N}}.$ (3)
The choice of classical reference algorithms is discussed in Sec. IV. Here, we focus specifically on the definition and measurement of runtime. The algorithm in Ref. [22] is a hybrid quantum–classical procedure consisting of two runs of SA and a single execution of the quantum routine. The first SA run provides an initial guess for the quantum step, and the output of the quantum routine is subsequently refined using a second SA run to mitigate potential readout errors.
The total runtime therefore includes both classical and quantum components. In Ref. [22], the classical runtime was estimated based on the time required to perform a single pass of SA, rather than being directly measured. In addition, overhead contributions associated with SA were not included, although these overheads are reported by the authors to be on the order of $1.6\,\mathrm{s}$. Including such contributions would affect the resulting runtime comparison. Similarly, the quantum runtime was estimated assuming a device throughput of $10^{4}$ circuit executions per second, without explicitly accounting for programming and transpilation overheads. A comprehensive definition of runtime for digital quantum devices should therefore include all of these processes.
Taken together, these considerations indicate that the reported runtime advantage in Ref. [22] is sensitive to the chosen definition and estimate of runtime. Under operationally grounded runtime metrics and consistent accounting of all relevant stages, a runtime advantage is not observed at the present stage. Even under the original measurement assumptions, comparison against stronger classical reference implementations alters the conclusions, as discussed in Sec. IV. A detailed analysis is provided in App. A3.
FIG. 1. : Time-to-epsilon $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ scaling with the size $N$ of the Sidon-28 instances, for values of $\varepsilon=0.75\%$ in panel (a), and $\varepsilon=1.25\%$ in panel (b). Results for QAC (blue) solver are reproduced from Ref. [20], courtesy of the authors. Remaining QAC data concerns experiments performed by us on the instances form Ref. [20]. With the same definition of runtime the results (orange) are consistent with  [20], which confirms correctness of methodology implementation. In addition $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ defined using the complete QPU access time $t_{f}^{{\rm QPU\,access}}$ as reported by D-Wave’s cloud interface (green), as well as runtime measured with cloud access $t_{f}^{{\rm runtime}}$ (red) are presented. Solid lines are power-law fits $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}\propto N^{\alpha}$, with the corresponding exponents $\alpha$ shown on the plot. The instance size range spanned by the lines denote which data points were used for the respective fits. Both plots clearly demonstrate that with the proper runtime definition $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ is constant for considered problem sizes. For comparison data for Simulated Bifurcation Machine (SBM) solver[37, 36, 39, 26] with $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ computed using 1 GPU and total runtime $t_{f}^{\rm tot}$ (cyan), as well as pure GPU runtime $t_{f}^{\rm GPU}$ (magenta) are presented. In this case the scaling of $\left[\mathrm{TT}\varepsilon\right]_{\mathrm{Med}}$ is more robust with respect to different runtime definitions, which is generally the case for classical algorithms. The bottom panel (c) shows detailed data of the fitting exponent and its uncertainty for values of $\varepsilon\in\{0.5,0.75,1.00,1.10,1.25,1.5\}\%$